OpenStudy (anonymous):

What is the value of x? If necessary, round your answer to two decimal places

OpenStudy (anonymous):



OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

nope

OpenStudy (skullpatrol):

Do you know the Pythagorean theorem: $$a^2+b^2=c^2$$

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

845?

OpenStudy (skullpatrol):



OpenStudy (anonymous):

no this not the answer

OpenStudy (anonymous):

13\[\sqrt{5}\]

OpenStudy (anonymous):

29.06

OpenStudy (anonymous):

no its 22.52 because u must rearrange the equation make a the subject and then take the square root of the and

OpenStudy (anonymous):

thanks